I need help on solving 3x + 4 = 10
Answer by doukungfoo(195)

First subtract 4 from both sides of the equation which results in
3x = 6
now to solve for x divde both sides of the equation by 3
3x + 4 = 10
3x + 4 - 4 = 10 - 4
3x = 6
3x/3 = 6/3
x = 2